a) An electron moving at 5 x 103 m/s in a 1.5T magnetic field experiences a magnetic force of 1.6× 10−16N. What angle does the velocity of the electron make with the magnetic field? 

b) A cosmic ray proton moving toward the earth at 5 x 107 m/s experiences a magnetic force of 1.70 × 10 −16N. What is the strength of the magnetic field if there is a 45° angle between it and the proton’s velocity? 




1
Expert's answer
2021-11-18T15:00:34-0500

Part(a)

We know that

"F=qvBsin\theta"

"\theta=sin^{-1}(\frac{F}{qvB})"

"\theta=sin^{-1}(\frac{1.6\times10^{-16}}{-1.6\times10^{-19}\times 5\times10^{3}\times1.5})"

"\theta=sin^{-1}(-0.13)=-7.66°"

Part (b)

"F=qvBsin\theta"

"B=\frac{F}{qvsin\theta}"


"B=\frac{1.7\times10^{-16}}{1.6\times10^{-19}\times5\times10^{7}\times sin45°}=3\times10^{-5}T"